OpenStudy (anonymous):

Solve this plz ..... 2log3(4) + 3log3(5)=10

hero (hero):

\[2\log_3(4) + 3\log_3(5) = 10\]

OpenStudy (lalaly):

\[\log_3(4^2)+\log_3(5^3)=10\]\[\log_3(16 \times 125)=10\]\[3^{10}=16 \times 125\]

OpenStudy (anonymous):

oh sorry if i replace 3 wid x then how will i solve it?

hero (hero):

nice la lady

OpenStudy (lalaly):

which 3?

hero (hero):

lalady, how's new zealand?

OpenStudy (lalaly):

its good, very cold :D but i love it, i go skiing everyday:D

OpenStudy (anonymous):

the base

hero (hero):

Oh wow. I hate cold and I hate skiing. But if someone like you escort me, I wouldn't care, lol

OpenStudy (lalaly):

haha!, let me know if u ever come here xD

hero (hero):

And how would I do that? I don't have any way of reaching you

OpenStudy (lalaly):

mmm... when u plan to visit, tell me here, then we'll c how u cn reach me :)

hero (hero):

Well, right before the financial bubble bursts, lol. Whenever that is

hero (hero):

Actually, if that happens, I may need to go somewhere warm. Like South America

OpenStudy (lalaly):

lols okay:D